On a recent quiz 6 students received course averaging 92 . The remaining students in the class received scores averaging 89 .if the quiz grades for the entire class averaged 90 .how many students were in the class ?

Answer by checkley77(12844)

6*92+89X=90(6+X)
552+89X=540+90X
89X-90X=540-552
-X=-12
X=12 STUDENTS AVERAGED 90 ON THE TESTS.
PROOF:
6*92+89*12=90(6+12)
552+1,068=90*18
1,620=1,620